OpenStudy (anonymous):

Solve 11.32x – 1 = 15.7 x ≈ 1.12 x ≈ 0.57 x ≈ 0.94 x ≈ 1.07 Yes I did get 1.47 as the answer but these are the 4 answers provided on the worksheet. Which one of these could it possibly be?

OpenStudy (jiteshmeghwal9):

11.32x=15.7+1 11.32x=16.7 x=16.7/11.32=?

OpenStudy (shaik0124):

the four option r wrong answer is1.475

OpenStudy (anonymous):

I know that's what I got. But THOSE are the 4 answers.

OpenStudy (shaik0124):

there might b mistake in d question